# Robust Amortized Bayesian Inference with Self-Consistency Losses on Unlabeled Data

### v4

@@ Acknowledgments
Daniel Habermann, Stefan Radev, and Paul Bürkner acknowledge support of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Projects 508399956 and 528702768. Paul Bürkner further acknowledges support of the DFG Collaborative Research Center 391 (Spatio-Temporal Statistics for the Transition of Energy and Transport) – 520388526.
@@ Appendix A Proofs
Appendix B Detailed setup of the multivariate normal case study
@@
The mean values $\mu_{\rm obs}\in\{0,1,\dots,11\}$ are progressively farther away from the training data. While conceptually simple and synthetic, this setting is already extremely challenging for simulation-based inference algorithms because of the large simulation gap [38]: standard NPEs are only trained on (labeled) training data that are several standard deviations away from the observed data the model sees at inference time.
The faithfulness of the approximated posteriors $q(\theta\mid x_{\rm obs})$ are assessed by computing the bias in posterior mean and standard deviation as well as the maximum mean discrepancy (MMD) with a Gaussian kernel [23] between the approximate and true (analytic) posterior.
@@
For the NPEs $q(\theta\mid x)$, we use a neural spline flow [8] with 5 coupling layers of 128 units each utilizing ReLU activation functions, L2 weight regularization with factor $\gamma=10^{-3}$, 5% dropout and a multivariate unit Gaussian latent space. The network is trained using the Adam optimizer for 100 epochs with a batch size of 32 and a learning rate of $5\times 10^{-4}$. These settings were the same for both the standard simulation-based loss and our proposed semi-supervised loss. For the conditions involving an estimated likelihood $q(x\mid\theta)$, we use the same configuration for the likelihood network as for the posterior network. For the summary network $h(x)$ (if included), we use a deep set architecture [45] with 30 summary dimensions and mean pooling, 2 equivariant layers each consisting of 2 dense layers with 64 units and a ReLU activation function. The inner and outer pooling functions also use 2 dense layers with the same configuration. The likelihood network as well as the summary network are jointly trained with the inference network using the Adam optimizer for 100 epochs with a batch size of 32 and a learning rate of $5\times 10^{-4}$.
Appendix C Comprehensive results for the multivariate normal case study
@@
Figure 6: Bias of posterior mean, bias of posterior standard deviation and posterior distance quantified by maximum mean discrepancy to the analytic posterior for variations of the default configuration outlined in Section 4.1. NPE approximators with the added self-consistency loss component are shown in blue, NPE approximators using just the standard simulation-based loss are shown in red. Irrespective of the varied factor and for all metrics, adding the self-consistency loss component is always a drastic improvement over the standard simulation-based loss alone. The plots show that adding the self-consistency loss component provides strong robustness gains even in high-dimensional spaces (top row) or when the self-consistency loss is evaluated on little data (center row). Variation of the mean of the unlabeled training data show that adding the self-consistency loss drastically improves posterior estimation as long as data used for evaluating the self-consistency loss is at least slightly out-of-distribution compared to the original training data ($\mu^{*}\geq 1$). Errorbars show ±$\pm 1$ standard deviations over 10 model refits on new training data.
@@
Figure 7: Bias of posterior mean, bias of posterior standard deviation and posterior distance quantified by maximum mean discrepancy to the analytic posterior when the likelihood is estimated (top row) and in presence of a summary network ($K=10$ data points; bottom row). In the setting where the likelihood function is estimated, we observe a lower bias of the posterior mean and lower maximum mean discrepancy to the true posterior when the self-consistency loss component is added compared to the standard simulation-based loss alone. However, we do see some bias of the posterior standard deviation, although with reversed signed compared to the standard loss. The self-consistency loss provides strong robustness gains in the presence of a summary network (and known likelihood) in terms of all metrics. Errorbars show ±$\pm 1$ standard deviations over 10 model refits on new training data.
Appendix D Detailed setup of the air traffic case study
@@
For the NPEs $q(\theta\mid x)$, we use a neural spline flow [8] with 6 coupling layers of 128 units each utilizing exponential linear unit activation functions, L2 weight regularization with factor $\gamma=10^{-3}$, 5% dropout and a multivariate unit Gaussian latent space. These settings were the same for both the standard simulation-based loss and our proposed semi-supervised loss. The simulation budget was set to $N=1024$. For the summary network, we use a long short-term memory layer with 64 output dimensions followed by two dense layers with output dimensions of 256 and 64. The inference and summary networks are jointly trained using the Adam optimizer for 100 epochs with a batch size of 32 and a learning rate of $5\times 10^{-4}$.
Appendix E Comprehensive results for the air traffic case study
Table 2: Posterior metrics for NPE and NPE augmented with self-consistency loss (NPE + SC) relative to Stan. For each parameter, the absolute bias in posterior means and standard deviations are reported along with the Wasserstein distance between the posteriors, using Stan as reference. The self-consistency loss was evaluated on data from $M=$ 4 countries during training. Metrics are averaged over all 15 countries.
Parameter $|\mu-\mu_{\rm Stan}|$ $|\sigma-\sigma_{\rm Stan}|$ Wasserstein distance
NPE NPE+SC NPE NPE+SC NPE NPE+SC
$\alpha$ $0.079$ 0.003 $0.033$ 0.020 $0.086$ 0.033
$\beta$ $0.153$ 0.012 $0.055$ 0.011 $0.161$ 0.070
$\gamma$ $0.087$ 0.048 $0.058$ 0.022 $0.154$ 0.112
$\delta$ $0.053$ 0.046 $0.038$ 0.018 $0.119$ 0.102
$\log(\sigma)$ $0.215$ 0.207 0.049 $0.058$ $0.304$ 0.282
Table 3: Posterior metrics for NPE and NPE augmented with self-consistency loss (NPE + SC) relative to Stan. For each parameter, the absolute bias in posterior means and standard deviations are reported along with the Wasserstein distance between the posteriors, using Stan as reference. The self-consistency loss was evaluated on data from $M=$ 15 countries during training. Metrics are averaged over all 15 countries.
Parameter $|\mu-\mu_{\rm Stan}|$ $|\sigma-\sigma_{\rm Stan}|$ Wasserstein distance
NPE NPE+SC NPE NPE+SC NPE NPE+SC
$\alpha$ $0.079$ 0.002 $0.033$ 0.002 $0.086$ 0.006
$\beta$ $0.153$ 0.001 $0.055$ 0.001 $0.161$ 0.009
$\gamma$ $0.087$ 0.002 $0.058$ 0.005 $0.154$ 0.014
$\delta$ $0.053$ 0.003 $0.038$ 0.005 $0.119$ 0.013
$\log(\sigma)$ $0.215$ 0.002 $0.049$ 0.004 $0.304$ 0.011
@@ Appendix E Comprehensive results for the air traffic case study
Parameter $\text{SE}(|\mu-\mu_{\rm Stan}|)$ $\text{SE}(|\sigma-\sigma_{\rm Stan}|)$
NPE SC(4) SC(8) SC(15) NPE SC(4) SC(8) SC(15)
$\alpha$ $0.019$ $0.009$ $0.013$ 0.001 $0.011$ $0.008$ $0.007$ 0.001
$\beta$ $0.024$ $0.026$ $0.023$ 0.002 $0.010$ $0.007$ $0.003$ 0.002
$\gamma$ $0.045$ $0.037$ $0.026$ 0.003 $0.015$ $0.016$ $0.013$ 0.003
$\delta$ $0.033$ $0.033$ $0.023$ 0.004 $0.015$ $0.016$ $0.012$ 0.002
$\log(\sigma)$ $0.076$ $0.076$ $0.058$ 0.002 $0.015$ $0.017$ $0.008$ 0.002
@@
Appendix F Detailed setup of the neuron activation case study
F.1 Model description

### v5

@@ Acknowledgments
Daniel Habermann, Stefan Radev, and Paul Bürkner acknowledge support of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Projects 508399956 and 528702768. Paul Bürkner further acknowledges support of the DFG Collaborative Research Center 391 (Spatio-Temporal Statistics for the Transition of Energy and Transport) – 520388526. Aayush Mishra acknowledges the computing time provided on the Linux HPC cluster at Technical University Dortmund (LiDO3).
Reproducibility Statement
We provide all details necessary to reproduce our results. Hyperparameter settings and network architectures are described in Section 4 and Appendices C, D, E, G, and I. The code to run all experiments is provided in the abstract. Hardware specifications are reported in Appendix K. All theoretical results are accompanied by complete proofs in Section 2 and Appendix A.
@@ Appendix A Proofs
Appendix B Frequently Asked Questions
Q: How do you define model-misspecification, simulation-gap, and domain shift?
We define model misspecification as a mismatch between an unknown (true) distribution of data $x^{*}\sim p^{*}(x)$ coming from a complex, real-world system, and the marginal likelihood of data $p(x|\mathcal{M})$ under some model $\mathcal{M}$. Concretely, we call a model misspecified if
$p^{*}(x)\neq p(x\mid\mathcal{M})\coloneqq\int_{\Theta}p(x\mid\theta,\mathcal{M})\,p(\theta\mid\mathcal{M})\,d\theta,\quad\forall x\in\mathcal{X}.$
Within this framework, a simulation gap refers to the divergence between the empirical distribution of observed data $p^{*}(x)$ and the empirical distribution of simulated data $\hat{p}(x)$, where $\hat{p}(x)$ is obtained via finite samples $x\sim p(x\mid\theta,\mathcal{M})$ for $\theta\sim p(\theta\mid\mathcal{M})$. That is, even when the model is not misspecified, limitations in simulation coverage can create gaps in support.
A domain shift refers to any distributional discrepancy between the simulated training data ($x\sim p(x\mid\mathcal{M})$) and the observed data ( $x^{*}\sim p^{*}(x)$). This includes, but is not limited to, model misspecification and simulation gaps, and encompasses broader generalization challenges faced when deploying amortized inference methods outside the simulation domain.
Q: What is the intuition behind using self-consistency to make the inference robust?
In a well-specified setting, the simulator accurately captures the true data-generating process and the learned posterior matches the true posterior. In this case, neural networks only suffer from approximation error due to finite data or model capacity. Under model misspecification, the observed data are drawn from a distribution $p^{*}(x)\neq p(x)$, and hence the model-implied posterior $p(\theta\mid x)\propto p(x\mid\theta)p(\theta)$ does not correspond to the posterior under $p^{*}(x)$ but under $p(x)$.
In such settings, neural networks face two types of error: approximation error, as before, due to imperfect learning, and extrapolation error, because the network encounters observations $x\sim p^{*}(x)$ that may lie outside the simulation support, and thus lacks guidance on how to sample from the posterior in those regions. In this case, minimizing the self-consistency loss encourages the learned posterior $q(\theta\mid x)$ to remain internally consistent with the assumed model $p(x,\theta)$ when evaluated on empirical samples $x\sim p^{*}(x)$. It enforces that the learned posterior $q(\theta\mid x)$ remains consistent with the model assumptions, effectively guiding the network to extrapolate in a model-coherent way. That is, even when the data are out-of-distribution relative to the simulator, the network learns to sample in a way that is self-consistent under the (misspecified) model posterior, improving robustness despite model mismatch.
Q: Why is the self-consistency loss coupled with standard simulation-based loss when self-consistency is strictly proper?
In principle, minimizing the self-consistency loss is sufficient to recover the true posterior, as guaranteed by Proposition 1. However, in practice, we observe that using the SC loss as the sole training objective leads to initialization and convergence issues. This is primarily due to the density-based nature of the SC loss. At initialization, when the approximate posterior and/or likelihood are poorly calibrated, the self-consistency loss can take on extreme values, often resulting in numerical instabilities, such as exploding gradients. Consequently, optimization tends to diverge or fail to make meaningful progress. We observe that the SC loss in tandem with the NPE loss provides a more stable and effective learning signal. The amortized loss anchors the model early on by using supervised simulated data, while the SC loss complements it by enforcing consistency on observed data, especially in out-of-simulation regimes.
Q: What is the motivation behind using the variance over the log Bayesian self-consistency ratio for $C$?
The choice of using the variance of the logarithm of the Bayesian self-consistency ratio, evaluated over $p\sim p_{C}(\theta)$ (Eq. 4), is motivated as follows: Köthe (2023) (Eq. 33) shows that a second-order Taylor expansion of the regular NPE loss function yields $KL(p(\theta\mid x)||q(\theta\mid x))\approx\frac{1}{2p(x)^{2}}\text{Var}_{p(\theta|x)}(\frac{p(\theta)p(x\mid\theta)}{q(\theta\mid x)})$. From this, it follows that minimizing $\text{Var}(\dots)$ also minimizes the KL divergence between true and approximate posterior. Because the denominator in $\text{Var}(\dots)$ can cause numerical instabilities due to vanishingly small values, we instead target $\text{Var}(\log(\dots))$. Minimizing this quantity via a loss function still minimizes the KL divergence between the true and approximate posterior, which follows from Proposition 1.
Q: What is the motivation behind using the approximate posterior $q_{t}(\theta\mid h(x^{*}))$ at training iteration $t$ for $p_{C}(\theta)$?
Rearranging Bayes’ rule shows that the marginal likelihood $p(x)=\frac{p(\theta)p(x\mid\theta)}{p(\theta\mid x)}$ is constant for all $\theta$. If we replace the true posterior $p(\theta\mid x)$ by an approximate posterior $q(\theta\mid x)$, large fluctuations in the marginal likelihood landscape are expected to be in high-density regions of the approximate posterior. We therefore choose the approximate posterior as the proposal distribution.
The choice of using the variance of the logarithm of the Bayesian self-consistency ratio, evaluated over $p\sim p_{C}(\theta)$ (Eq. 4), is motivated as follows: Köthe (2023) (Eq. 33) shows that a second-order Taylor expansion of the regular NPE loss function yields $KL(p(\theta\mid x)||q(\theta\mid x))\approx\frac{1}{2p(x)^{2}}\text{Var}_{p(\theta|x)}(\frac{p(\theta)p(x\mid\theta)}{q(\theta\mid x)})$. From this, it follows that minimizing $\text{Var}(\dots)$ also minimizes the KL divergence between true and approximate posterior. Because the denominator in $\text{Var}(\dots)$ can cause numerical instabilities due to vanishingly small values, we instead target $\text{Var}(\log(\dots))$. Minimizing this quantity via a loss function still minimizes the KL divergence between the true and approximate posterior, which follows from Proposition 1.
Q: How to choose the self-consistency weight $\lambda$?
From the strict-properness of self-consistency losses, it follows that $\lambda$ does not control a trade-off between the standard NPE loss and the SC loss, as both optimize for the same training objective. Instead, its main role is to stabilize training by preventing initialization and convergence issues. For experiments with low-dimensional parameter spaces, a default choice of $\lambda=1$ typically works well. However in high-dimensional settings, such as the image-denoising experiment, the approximate posterior can be particularly bad in the early epochs, leading to unstable training. In such cases, we found it to be effective to use a linear schedule for $\lambda$ that keeps it at 0 for the first few epochs and then gradually increases to 1.
Appendix C Detailed setup of the multivariate normal case study
@@
The mean values $\mu_{\rm obs}\in\{0,1,\dots,11\}$ are progressively farther away from the training data. While conceptually simple and synthetic, this setting is already extremely challenging for simulation-based inference algorithms because of the large simulation gap (Schmitt et al., 2023): standard NPEs are only trained on (labeled) training data that are several standard deviations away from the observed data the model sees at inference time.
The faithfulness of the approximated posteriors $q(\theta\mid x_{\rm obs})$ are assessed by computing the bias in posterior mean and standard deviation as well as the maximum mean discrepancy (MMD) with a Gaussian kernel (Gretton et al., 2012) between the approximate and true (analytic) posterior.
@@
For the NPEs $q(\theta\mid x)$, we use a neural spline flow (Durkan et al., 2019) with 5 coupling layers of 128 units each utilizing ReLU activation functions, L2 weight regularization with factor $\gamma=10^{-3}$, 5% dropout and a multivariate unit Gaussian latent space. The network is trained using the Adam optimizer for 100 epochs with a batch size of 32 and a learning rate of $5\times 10^{-4}$. These settings were the same for both the standard simulation-based loss and our proposed semi-supervised loss. For the conditions involving an estimated likelihood $q(x\mid\theta)$, we use the same configuration for the likelihood network as for the posterior network. For the summary network $h(x)$ (if included), we use a deep set architecture (Zaheer et al., 2017) with 30 summary dimensions and mean pooling, 2 equivariant layers each consisting of 2 dense layers with 64 units and a ReLU activation function. The inner and outer pooling functions also use 2 dense layers with the same configuration. The likelihood network as well as the summary network are jointly trained with the inference network using the Adam optimizer for 100 epochs with a batch size of 32 and a learning rate of $5\times 10^{-4}$.
Appendix D Comprehensive results for the multivariate normal case study
@@
Figure 6: Bias of posterior mean, bias of posterior standard deviation and posterior distance quantified by maximum mean discrepancy to the analytic posterior for variations of the default configuration outlined in Section 4.1. NPE approximators with the added self-consistency loss component are shown in blue, NPE approximators using just the standard simulation-based loss are shown in red. Irrespective of the varied factor and for all metrics, adding the self-consistency loss component is always a drastic improvement over the standard simulation-based loss alone. The plots show that adding the self-consistency loss component provides strong robustness gains even in high-dimensional spaces (top row) or when the self-consistency loss is evaluated on little data (center row). Variation of the mean of the unlabeled training data show that adding the self-consistency loss drastically improves posterior estimation as long as data used for evaluating the self-consistency loss is at least slightly out-of-distribution compared to the original training data ($\mu^{*}\geq 1$). Errorbars show $\pm 1$ standard deviations over 10 model refits on new training data.
@@
Figure 7: Bias of posterior mean, bias of posterior standard deviation and posterior distance quantified by maximum mean discrepancy to the analytic posterior when the likelihood is estimated (top row) and in presence of a summary network ($K=10$ data points; bottom row). In the setting where the likelihood function is estimated, we observe a lower bias of the posterior mean and lower maximum mean discrepancy to the true posterior when the self-consistency loss component is added compared to the standard simulation-based loss alone. However, we do see some bias of the posterior standard deviation, although with reversed signed compared to the standard loss. The self-consistency loss provides strong robustness gains in the presence of a summary network (and known likelihood) in terms of all metrics. Errorbars show $\pm 1$ standard deviations over 10 model refits on new training data.
D.1 Comparison to other robustness methods
Existing approaches for enhancing the robustness of amortized Bayesian inference typically fall into one of three categories, each with its own limitations. First, regularization-based methods (e.g., Gloeckler et al., 2023; Huang et al., 2023) introduce additional constraints inevitably altering the original training objective, thereby deviating from the true Bayesian posterior that the self-consistency loss seeks to approximate.
Second, some methods employ supervised corrections using a small amount of real-world, labeled data (e.g., Wehenkel et al., 2024) which requires access to ground-truth parameters, shifting the setting from semi-supervised to fully supervised learning.
Third, unsupervised domain adaptation (UDA) techniques such as NPE-DANN and NPE-MMD have been proposed to improve robustness under domain shift by aligning the distributions of simulated and observed data in the summary feature space (Elsemüller et al., 2025). These losses govern the alignment of the summary space between simulated and observed data, effectively adjusting the observed data resulting in posteriors conditioned on a transformed version of the observation (Target 2, Section 3, Elsemüller et al., 2025), rather than the analytic posterior. Moreover, as shown by Elsemüller et al. (2025), these UDA methods and NNPE (Ward et al., 2022) exhibit notably poor performance under prior misspecification.
That said, we conducted additional experiments with NPE-MMD (Huang et al., 2023), NPE-DANN (Elsemüller et al., 2025), and NNPE (Ward et al., 2022) on the two-dimensional multivariate normal model, replicating the prior misspecification setting described in Section 4.1. Since both UDA methods rely on a summary network for feature alignment, we employed a shared DeepSet architecture (2-layer network with 30-dimensional output) across all methods, including NPE+SC, to ensure a fair comparison. For self-consistency loss, we used 32 unlabeled observed data at $\mu_{obs}=3$ during training. We used Adam optimizer with a learning rate of 0.0005 trained for 40 epochs using 1,028 simulations. All the hyperparameter configurations were kept the same as described in Section 4.1 and Appendix C. For NPE-MMD, NPE-DANN, and NNPE, we utilized the implementation by Elsemüller et al. (2025).
The results clearly show that NPE-MMD, NPE-DANN, and NNPE fail under prior misspecification. Also, with increase in misspecification, the results for these methods became even worse while NPE+SC maintains the robustness gains. UDA methods also require large unlabeled data during training while we see strong robustness gains in NPE+SC with as few as 4 unlabeled data points. As these methods have a different training objective (Target 2, Section 3, Elsemüller et al., 2025) rather than the analytic posterior, we do not include them in other case study comparisons. Instead, we focus on standard simulation-based NPE and state-of-the-art MCMC via Stan, which target the analytic posterior and serve as appropriate baseline and benchmark, respectively, for evaluating our semi-supervised NPE+SC approach.
Figure 8: Bias of posterior mean, bias of standard deviation, and posterior distance quantified by maximum mean discrepancy, relative to the analytic posterior. NPE + SC provides far more robustness gains compared to NPE-MMD, NPE-DANN, and NNPE for the case of prior misspecification. Errorbars show $\pm 1$ standard deviations over 10 model refits on new training data.
Appendix E Detailed setup of the air traffic case study
The parameters $\alpha_{j}$ are country-level intercepts, $\beta_{j}$ are the autoregressive coefficients, $\gamma_{j}$ are the regression coefficients of household debt and $\delta_{j}$ are the regression coefficients of GDP per capita, and $\sigma_{j}$ is the standard deviation of the noise term. This model was previously used within ABI in Habermann et al. (2024). As commonly done for autoregressive models, we regress on time period differences to mitigate non-stationarity. This is critical for simulation-based inference because when $\beta_{j}>1$, exponential growth quickly produces unrealistic air traffic volumes. Moreover, amortizing over covariate spaces, such as varying GDP per capita between countries, can lead to model misspecification if such fluctuations are underrepresented in training.
@@ Appendix E Detailed setup of the air traffic case study
For the NPEs $q(\theta\mid x)$, we use a neural spline flow (Durkan et al., 2019) with 6 coupling layers of 128 units each utilizing exponential linear unit activation functions, L2 weight regularization with factor $\gamma=10^{-3}$, 5% dropout and a multivariate unit Gaussian latent space. These settings were the same for both the standard simulation-based loss and our proposed semi-supervised loss. The simulation budget was set to $N=1024$. For the summary network, we use a long short-term memory layer with 64 output dimensions followed by two dense layers with output dimensions of 256 and 64. The inference and summary networks are jointly trained using the Adam optimizer for 100 epochs with a batch size of 32 and a learning rate of $5\times 10^{-4}$.
Appendix F Comprehensive results for the air traffic case study
Table 1: Posterior metrics for NPE and NPE augmented with self-consistency loss (NPE+SC) relative to Stan. For each parameter, absolute bias in posterior means and standard deviations, as well as Wasserstein distance, are reported. Values are shown as mean (SE) across 15 countries. Self-consistency loss was evaluated using $M\in\{4,8,15\}$ countries during training.
Parameter NPE NPE+SC ($M=4$) NPE+SC ($M=8$) NPE+SC ($M=15$)
Mean bias ($|\mu-\mu_{\rm Stan}|$)
$\alpha$ $079\pm 0.019 0.0$$03\pm 0.009 0.0$$14\pm 0.013 0.0$$002\pm 0.001$
$\beta$ $153\pm 0.024 0.0$$12\pm 0.026 0.0$$31\pm 0.023 0.0$$001\pm 0.002$
$\gamma$ $087\pm 0.045 0.0$$48\pm 0.037 0.0$$06\pm 0.026 0.0$$002\pm 0.003$
$\delta$ $053\pm 0.033 0.0$$46\pm 0.033 0.0$$42\pm 0.023 0.0$$003\pm 0.004$
$\log(\sigma)$ $0.215\pm 0.076$ $0.207\pm 0.076$ $148\pm 0.058 0.0$$002\pm 0.002$
Standard deviation bias ($|\sigma-\sigma_{\rm Stan}|$)
$\alpha$ $033\pm 0.011 0.0$$20\pm 0.008 0.0$$20\pm 0.007 0.0$$002\pm 0.001$
$\beta$ $055\pm 0.010 0.0$$11\pm 0.007 0.0$$04\pm 0.003 0.0$$001\pm 0.002$
$\gamma$ $058\pm 0.015 0.0$$22\pm 0.016 0.0$$35\pm 0.013 0.0$$005\pm 0.003$
$\delta$ $038\pm 0.015 0.0$$18\pm 0.016 0.0$$31\pm 0.012 0.0$$005\pm 0.002$
$\log(\sigma)$ $049\pm 0.015 0.0$$58\pm 0.017 0.0$$11\pm 0.008 0.0$$004\pm 0.002$
Wasserstein distance
$\alpha$ $086\pm 0.019 0.0$$33\pm 0.009 0.0$$35\pm 0.013 0.0$$006\pm 0.001$
$\beta$ $161\pm 0.024 0.0$$70\pm 0.026 0.0$$54\pm 0.023 0.0$$009\pm 0.002$
$\gamma$ $0.154\pm 0.045$ $112\pm 0.037 0.0$$68\pm 0.026 0.0$$014\pm 0.003$
$\delta$ $0.119\pm 0.033$ $102\pm 0.033 0.0$$64\pm 0.023 0.0$$013\pm 0.002$
@@ Appendix F Comprehensive results for the air traffic case study
F.1 Comparison to VAE and online training of NPE
We conducted additional experiments using a variational auto-encoder (VAE) (Kingma and Welling, 2013) baseline with a known decoder (likelihood) and a neural encoder (posterior network). The encoder is a feed-forward network with three hidden layers of 128 units each, followed by two output heads that parameterize the mean and standard deviation of a standard normal base distribution. The model was trained on 40,960 simulated pairs until convergence. We trained NPE and NPE + SC in online setting that leads to training data of the same size as VAE. We also compare our method against Schmitt et al. (2024) where simulated data is used to evaluate SC loss during training. We similarly used 40,960 simulations for training NPE and 32 simulations for evaluating self-consistency loss for 40 epochs.
Results comparing the above methods for the air-traffic experiment are reported in Table 2 and Figure 12. These results show that the performance of VAE with known likelihood is comparable to NPE (which does not use explicit likelihood density) but it is much worse than our semi-supervised (NPE + SC) approach. Our results also show that the method of Schmitt et al. (2024) does not overcome model misspecification and performs poorly compared to our semi-supervised approach.
Table 2: Posterior Wasserstein distances, mean bias, and standard deviation bias for online trained NPE, Schmitt et al. (2024), VAE, and our proposed NPE + SC method relative to Stan. The self-consistency loss was evaluated on data from $M=\textbf{15}$ countries during training. Metrics are averaged over all 15 countries.
Parameter Mean bias ($|\mu-\mu_{\rm Stan}|$)
NPE (online) Schmitt et al. (2024) VAE NPE (online) + SC
$\alpha$ $075\pm 0.028 0.0$$26\pm 0.019 0.0$$60\pm 0.017 0.0$$008\pm 0.002$
$\beta$ $087\pm 0.020 0.0$$52\pm 0.028 0.0$$96\pm 0.020 0.0$$013\pm 0.003$
$\gamma$ $0.101\pm 0.014$ $0.200\pm 0.071$ $108\pm 0.020 0.0$$014\pm 0.003$
$\delta$ $0.096\pm 0.018$ $12\pm 0.072 0.110$$10\pm 0.028$ $0.009\pm 0.002$
$\log(\sigma)$ $189\pm 0.051 0.0$$042\pm 0.141$ $142\pm 0.024 0.0$$009\pm 0.002$
Standard deviation bias ($|\sigma-\sigma_{\rm Stan}|$)
$\alpha$ $030\pm 0.008 0.0$$03\pm 0.021 0.0$$29\pm 0.006 0.0$$005\pm 0.001$
$\beta$ $011\pm 0.002 0.0$$39\pm 0.039 0.0$$50\pm 0.004 0.0$$011\pm 0.002$
$\gamma$ $058\pm 0.016 0.0$$22\pm 0.033 0.0$$79\pm 0.014 0.0$$009\pm 0.002$
$\delta$ $059\pm 0.017 0.0$$01\pm 0.041 0.0$$83\pm 0.015 0.0$$016\pm 0.003$
$\log(\sigma)$ $170\pm 0.009 0.0$$66\pm 0.034 0.0$$26\pm 0.006 0.0$$013\pm 0.002$
Wasserstein distance
$\alpha$ $080\pm 0.028 0.0$$68\pm 0.017 0.0$$65\pm 0.016 0.0$$010\pm 0.002$
$\beta$ $0.088\pm 0.020$ $0.133\pm 0.024$ $102\pm 0.019 0.0$$018\pm 0.002$
$\gamma$ $0.113\pm 0.015$ $0.278\pm 0.050$ $125\pm 0.019 0.0$$018\pm 0.003$
$\delta$ $0.110\pm 0.020$ $0.240\pm 0.049$ $128\pm 0.027 0.0$$017\pm 0.002$
$\log(\sigma)$ $0.227\pm 0.045$ $0.459\pm 0.073$ $144\pm 0.024 0.0$$016\pm 0.002$
Figure 12: Comparison of posterior estimates among online trained NPE, Schmitt et al. (2024), VAE, and our proposed semi-supervised NPE+SC method against Stan (reference). The self-consistency loss was evaluated on data from $M=$ 15 countries during training. The results clearly show that our method is consistent with Stan and outperforms all other methods by a large margin.
Appendix G Detailed setup of the neuron activation case study
G.1 Model description